A body is describing a circular motion with a speed v when it describes an angle α at the centre the change in the velocity delta v is
1
Expert's answer
2020-03-31T09:24:14-0400
"v_1=v,v_2=v"

"\Delta v=|\bold{v_2-v_1}|=\sqrt{v_1^2+v_2^2-2v_1v_2\cos\alpha}"

"\Delta v=\sqrt{2v^2(1-\cos\alpha)}"

"\Delta v=2v\sin{\frac{\alpha}{2}}"